OpenStudy (anonymous):

Can someone help me? Solve. 5^2x-5=125

OpenStudy (jdoe0001):

\(\large \bf 5^{2x-5}=125 ?\)

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

well.. any ideas on what's say \(\large 5^3=?\)

OpenStudy (anonymous):

125?

OpenStudy (jdoe0001):

so, we could say that \(\bf 5^{2x-5}=125\implies 5^{{\color{brown}{ 2x-5 }}}=5^{\color{brown}{ 3 }} \\ \quad \\ \textit{that means that }\qquad 2x-5 = 3\)

OpenStudy (jdoe0001):

same base on both sides the exponents "must" be the same as well thus they equal each other